OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation 4(2x − 3) = 20 is given below. Choose the step that is missing from the solution. 4(2x − 3) = 20 8x = 32 x = 4 6x − 3 = 24 8x − 12 = 20 2x − 12 = 24 6x + 12 = 5

OpenStudy (cwrw238):

missing step comes after 4(2x - 3) = 20 distribute 4 over the parentheses will give the missing step

OpenStudy (cwrw238):

4( a - b) = 4a - 4b

OpenStudy (anonymous):

is it c?

OpenStudy (cwrw238):

no what is 4 * 2x ?

OpenStudy (anonymous):

8 so its b

OpenStudy (cwrw238):

right!

OpenStudy (cwrw238):

4(2x - 3) = 8x - 12